OpenStudy (amenah8):

How would you solve this absolute value equation? l f + 10 l = 1 Also, how do you solve absolute value equations in general?

jimthompson5910 (jim_thompson5910):

Rule: if |x| = k, then x = k or x = -k where k is some positive number

OpenStudy (anonymous):

subtract 10 from both sides

jimthompson5910 (jim_thompson5910):

in this particular case, `l f + 10 l = 1` breaks down into `f+10=1 or f+10=-1`

OpenStudy (anonymous):

f=-9 or f=-11

OpenStudy (amenah8):

Alright, thanks!